# Predicting potentially unfair clauses in Chilean terms of services with natural language processing

### v1

@@ 3.5 Gray Clauses
Gray Clauses are provisions that, in contrast with the Dark Clauses, are not easily identifiable as unfair. Their legal base is Article 16 letter g) LPC, that refers, in general, to clauses that are contrary to "good faith". That is, they cause a significant imbalance in the rights and obligations of the parties of the contract in detriment of the consumers. This definition is broad and overlaps with the concept of the unfair clause itself. It also does not provide clear guidelines to proceed. On the contrary, it is meant to constitute an open category to include any potentially unfair clauses not included by the legislator in the previous cases of the norm. The specific situations comprised under this umbrella provision are a matter of interpretation. Due to this there is no presumption of unfairness. That means, that the consumer will have to prove in court that the disputed clause is unfair under the terms of Article 16 letter g) LPC  [4].
@@
The ambiguity of this category makes it the hardest of all the ones analyzed, both for the consumers, the legal experts and also for ML models. Their identification will usually depend on the interconnected understanding of different sections of the ToS and even of contextual facts not present in the wording of the documents. Furthermore, their determination will also depend on the legal argumentation of the plaintiff and defendants, and finally, on the opinion of the court.
@@
However, during the analysis of the ToS, our legal experts were able to identify eight different classes of Gray Clauses, based on the contractual imbalance that they generate, and on the frequency that they were found. Examples of these clauses are the ones that impose risks or cost on the consumer arising from external factors outside of the contractual relationship, such as force majeur  [^4] This refers to unforeseeable circumstances that are out of the control of the parties of a contract, for example, natural disasters; that grant the provider the ability to share information of the consumer with unrelated parties, or that grant the provider broad, perpetual, or irrevocable faculties that extend beyond the contract’s duration; among other clauses. These classes of Gray Clauses are suggestions of our legal experts, but could be extended by analyzing other types of consumer contracts and ToS, for example, of financial services.
@@
Whether a clause is defined as Dark or Gray may also be a matter of degree that depends on the legal argumentation of the reviewer. For example, Article 16 letter a) LPC clearly states that clauses that allow the modification of contracts at the sole discretion of the provider or allow its unilateral termination are not permitted. These would be considered as Dark Clauses under our categorization, as it was the case when we found clauses that allow changes or terminations of services without previous notice during the revision of the ToS. However when a change was notified with enough anticipation, for example, a month in advance, the legal experts considered that the "unilateral" nature of the change was altered, since the consumer may have the opportunity to be aware of the future change and possibly negotiate the modifications or its consequences. This might be only potentially feasible, since the consumer might still not have any bargaining power to prevent the change and the only alternative remaining might be to leave the service. The clause shows an imbalance between the parties in detriment of the consumer and therefore was label as a Gray Clause, but since the outcome is equivalent to the first example, this one could also be considered as Dark Clauses by an stricter interpreter.
@@
For the purpose of the annotations and the learning process of the model, we considered it very valuable to also distinguish these slightly different variations of similar circumstances. As in the example of unilateral modification, a model trained on samples without previous notification by the provider may have low precision once the grammatical structure of the sample is changed to include contractual changes with previous notifications. For consumers it is valuable to highlight different potentially unfair practices to raise the awareness to these situations.
@@ 3.6 Annotated Clauses
From the four categories of Unfair Clauses identified in our research, we analyze three with our ML methodology: the Illegal, the Dark and the Gray Clauses. For these categories we select 20 classes according to the legal source used to define them and the frequency they appear in ToS. This section describes the annotated clauses in detail, introducing their keys in the dataset, their legal sources and their explanations in the following Table 1.
@@
lpc int Clauses that modify the liability of intermediaries of good or services Art. 43 LPC
@@
lpc Clauses that violate a provision of the LPC that does not refer to the list of unfair clauses of Article 16. This constitutes a residual class LPC
@@
na Clauses that mention a foreign law, which are not applicable in Chile Art. 16 CC
@@
cr Clauses that include an unilateral and arbitrary modification of the ToS without the consumer being able to challenge or reject the change Art. 16 a) LPC
ter Clauses that include an unilateral and arbitrary termination of services Art. 16 a) LPC
@@
des risk Clauses that imposes risks or costs on the consumer that are unrelated to the provider arising from external factors to the contractual relationship, such as force majeure Art. 16 g) LPC
des def Clauses that impose obligations on the consumer to defend, indemnify or cooperate in potential litigation in favor of the provider Art. 16 g) LPC
@@
bfe Clauses that go against the requirements of good faith taking in consideration the purpose of the contract. This constitutes a residual class Art. 16 g) LPC
Table 1: The annotation scheme consists of the potentially unfair clauses’ keys in the dataset, their legal sources and their explanations
@@ 3.7 Corpus
The 50 ToS analyzed comprise of 6,523 clauses. Two legal experts applied the annotation scheme on these clauses and marked them with one or more keys, refer to Table 1 for details. Of the 20% that are identified as abusive, around 45% are categorized as illegal, 26% as dark and 28% as gray. For a detailed breakdown, refer to Table 8. Following Section 3.6, we aggregate and recombine the annotated clauses into the datasets Illegal, Dark and Gray Clauses. For each category, we define the binary classification task of detection and the multi-label multi-class task of classification. We eliminate any clause that comprises fewer than seven words. This yields a total corpus size of 15,201 annotated clauses over six datasets.
For the detection task, we reduce the labels for the potentially unfair clauses to the uniform label "unfair" and and other clauses to the label "ok", see Table. 2 for an overview. The classification task then consists of only the paragraphs that are labeled potentially unfair, see Table. 3 for an overview. These tasks are inherently multi-label, and it is common to find combinations of labels for a single clause, see the co-occurrence in Fig. 1. However, our corpus exhibits varying levels of co-occurrence, ranging from the almost pure annotations in Illegal Classification to the high rates of co-occurrence in Dark Classification and particularly in Gray Classification.
@@
Since some combinations of multi-label annotations are extremely rare, we employed an iterative stratified sampling method [58] to account for the significant disparity in the number of distinct label sets. This method ensures a highly representative split into training, validation and test, given the complexity of the annotations. Our use of this method involves dividing each subset into training and testing sets by a ratio of 70% to 30%, and then further subdividing the training set into training and validation sets, with a ratio of 70% to 30%. Due to the multi-label nature of our classification datasets, the sum of instances per label after stratified sampling may be different from the detection task. For example, the algorithm selectively over-samples in order to stratify the Illegal Classification dataset split to a total of 643 samples.
@@ 4 Methodology
We employ two learning paradigms to detect and to classify potentially unfair clauses. First, we fine-tune transformer-based models. Second, we evaluate the few-shot performance of commercial and open LLMs. We follow [13] and include a linear Support Vector Machine [16] as a baseline. This SVM uses TF-IDF features for the top-$K$ frequent $n$-grams ($n\in[1,2,3]$).
@@ 4.1 Fine-Tuning Classifier Models
  • BERT [18] is an encoder-model pre-trained on a corpus of English texts. It performs masked language modeling with a vocabulary of 32K tokens and up to a sequence length of 512 tokens. Longer texts are usually truncated. We include it as an essential baseline method.
  • Multilingual BERT (M-BERT) [18] is trained on 104 languages, but otherwise, it’s the same as BERT. We include it in our study because M-BERT generates multilingual representations that are easy to fine-tune and achieve excellent performance in Spanish-language tasks [50].
@@
For the detection tasks we use a binary cross entropy loss. For the multi-label classification task with $n$ classes, we use the binary cross entropy loss with logits, where a binary vector of length $n$ encodes the possible classes. We report the micro-F1 and macro-F1 scores for all models and datasets since the classes are highly imbalanced. Besides these, we report precision and recall metrics for detailed inspection. To calculate the multi-label metrics, we convert the networks’ sigmoid output signal to binary predictions for the $n$ classes [13].
@@ 4.2 Prompting Large Language Models
The evaluation of the few-shot capabilities of Large Language Models focuses on models that explicitly support multilingual data and have sufficiently long context lengths to process the constructed, long few-shot queries. This excludes models that can only process fewer than 8,000 tokens at a time. We categorize the selected LLMs into commercial or open models. For an overview, refer to Table 4.
@@
We prompt the LLMs using the following structure. A header describes the task and provides a list with possible labels to choose from:
    Dados los siguientes ejemplos de clusulas chilenas y sus
    etiquetas, prediga la ltima clusula segn su etiqueta.
    Responda slo con las etiquetas correctas. No explique su
    decisin. A continuacin, se indican todas las opciones
    posibles, seguidas de ejemplos tpicos.
This translates into the following task description: "Given the following examples of clauses and their labels, predict the last clause by their label. Respond with the correct labels only. Do not explain your decision. Here are all possible options, followed by typical examples." This is then followed by one of the two following sentences, depending on whether the task is classification or detection.
    Una clusula puede tener entre una y tres etiquetas.
    Una clusula slo puede tener una etiqueta.
These translate to either "A clause can have between one and three labels" or "A clause can have only one label." This is followed by a list of possible clauses to predict. For instance, when classifying potentially illegal clauses, we add the following list:
@@
    Label:
@@ 5 Experiments
Experimental setup. We obtain pre-trained models from either the Huggingface Hub  [^5] https://huggingface.co/models, transformers version 4.44.0 or the Ollama Library  [^6] https://ollama.com/library, Ollama version 0.3.11. For few-shot learning, we select "instruct" variants that are tuned to follow prompts. We use 8-bit quantization, because this level of quantization increases throughput considerably, while still recommended for processing long texts because its negative effect on performance is negligible [38]. We construct the few-shot prompts from the combined training and validation datasets. For each model we perform runs with three different random seeds and report their average test set scores with standard deviation. We fine-tune the smaller models using a binary cross entropy loss with multi-label logits, and use the Adam optimizer [35] with a learning rate of 3e-5 and a batch size of 32 with up to 50 epochs. Where necessary, we apply gradient accumulation to obtain the target batch size. We regularize the training further with an early-stopping mechanism set to 10 epochs of patience. It tracks the macro-F1 score of the validation split because the datasets are imbalanced. Our fine-tuning experiments extend the code base published by LexGlue [13] [^7] https://github.com/coastalcph/lex-glue, commit 20ad76a and accelerate training using mixed precision on NVIDIA Ada Lovelace GPUs.
Since the datasets are highly imbalanced we do not only report the micro-F1 score ($\mu$-F1), which provides a global measure but may be dominated by the majority classes. We also report the macro-F1 (m-F1) score, that allows us to evaluate models on the classes of each dataset equally, highlighting the performance on minority classes. For the best-performing fine-tuned models we additionally break down the results into precision and recall for each classification task, see Sec. 5.3. Our study aggregates the results from a total of 426 experiments.
@@ 5.1 Detection
We demonstrate the ability of LLMs to detect potentially abusive clauses in the Chilean ToS corpus. To this end, we conducted large-scale experiments with a diverse set of models on the three detection datasets of the corpus and report the $\mu$-F1 and m-F1 scores and their standard deviation per model over three runs. The results in Tab. 5 are subdivided in four groups of models: i) an SVM baseline that was trained from scratch, ii) six transformer models pre-trained on English, iii) eight transformer models pre-trained with Spanish or multiple languages, and iv) four LLMs queried via few-shot learning.
@@
Table 5: Experimental results for the three detection datasets. For convenience, we print the overall best results in bold, and underline the results that are best per group. The results show that detection is feasible. Moreover, they indicate a difficulty-ranking of the corpora with the Illegal Detection dataset on the one hand being easier to predict and Dark Detection and Gray Detection on the other hand harder, resulting in lower detection rates.
@@
On the Dark Detection dataset, the baseline SVM delivers the highest $\mu$-F1 Score of 96% and BETO uncased the highest m-F1 Score of 80%. The Dark Clauses are harder to detect than Illegal Clauses. This may be due to their relative lower occurrence in the corpus and their imbalance, leading to lower m-F1 scores. Most English transformers exhibit a very high variance in m-F1 scores and significant drops in performance. The regularization may not completely eliminate overfitting to more frequent classes, possibly leading to this problematic higher variance. However, more appropriate pre-training remedies these problems as multilingual or Spanish-language pre-training demonstrates. RoBERTalex performs on par with M-BERT cased and only falls behind BETO and M-BERT uncased by 2% m-F1 Score. Finally, few-shot learning is ineffective and the best models are behind by 10% in $\mu$-F1 and 14% $m$-F1.
@@
In summary, our proposed detection datasets present their own unique challenges and our analysis emphasizes the importance of selecting a suitable model architecture and dataset characteristics, such as pre-training corpora. However, the strong performance of the SVM and of small, appropriately pre-trained transformers demonstrates, that the detection of potentially abusive clauses in the Chilean ToS corpus is possible, with the rates of $96\%$ $\mu$-F1 and $89\%$ m-F1 on Illegal Detection, $96\%$ $\mu$-F1 and $80\%$ m-F1 on Dark Detection, $96\%$ $\mu$-F1 and $79\%$ m-F1 on Gray Detection. This allows us to separate the detection from subsequent classification, which we will discuss in the next chapter.

### v2

@@ 3.5 Gray Clauses
Gray Clauses are provisions that, in contrast with the Dark Clauses, are not easily identifiable as abusive. Their legal base is Article 16 letter g) LPC, which refers, in general, to clauses that are contrary to "good faith". That is, they cause a significant imbalance in the rights and obligations of the parties of the contract to the detriment of the consumers. This definition is broad and overlaps with the concept of the abusive clause itself. It also does not provide clear guidelines to proceed. On the contrary, it is meant to constitute an open category to include any potentially abusive clauses not accounted for by the legislator in the previous cases of the norm. The specific situations comprised under this umbrella provision are a matter of interpretation. Due to this, there is no presumption of abusiveness. That means, that the consumer will have to prove in court that the disputed clause is abusive under the terms of Article 16 letter g) LPC [4].
@@
The ambiguity of this category makes it the hardest of all the ones analyzed, both for the consumers, the legal experts, and also for ML models. Their identification will usually depend on the interconnected understanding of different sections of the ToS and even of contextual facts not present in the wording of the documents. Furthermore, their determination will also depend on the legal argumentation of the plaintiff and defendants, and finally, on the opinion of the court.
@@
However, during the analysis of the ToS, our legal experts were able to identify eight different classes of Gray Clauses, based on the contractual imbalance that they generate, and on the frequency that they were found. Examples of these clauses are the ones that impose risks or costs on the consumer arising from external factors outside of the contractual relationship, such as force majeur [^4] This refers to unforeseeable circumstances that are out of the control of the parties of a contract, for example, natural disasters; that grant the provider the ability to share information of the consumer with unrelated parties, or that grant the provider broad, perpetual, or irrevocable faculties that extend beyond the contract’s duration; among other clauses. These classes of Gray Clauses are suggestions of our legal experts but could be extended by analyzing other types of consumer contracts and ToS, for example, financial services.
@@
Whether a clause is defined as Dark or Gray may also be a matter of degree that depends on the legal argumentation of the reviewer. For example, Article 16 letter a) LPC clearly states that clauses that allow the modification of contracts at the sole discretion of the provider or allow its unilateral termination are not permitted. These would be considered Dark Clauses under our categorization, as was the case when we found clauses that allow changes or terminations of services without previous notice during the revision of the ToS. However, when a change was notified with enough anticipation, for example, a month in advance, the legal experts considered that the "unilateral" nature of the change was altered, since the consumer may have the opportunity to be aware of the future change and possibly negotiate the modifications or its consequences. This might be only potentially feasible since the consumer might still not have any bargaining power to prevent the change and the only alternative remaining might be to leave the service. The clause shows an imbalance between the parties to the detriment of the consumer and therefore was labeled as a Gray Clause, but since the outcome is equivalent to the first example, this one could also be considered a Dark Clause by a stricter interpreter.
@@
For the annotations and the learning process of the model, we considered it very valuable to also distinguish these slightly different variations of similar circumstances. As in the example of unilateral modification, a model trained on samples without previous notification by the provider may have low precision once the grammatical structure of the sample is changed to include contractual changes with previous notifications. For consumers, it is valuable to highlight different potentially abusive practices to raise awareness of these situations.
@@ 3.6 Annotated Clauses
From the four categories of Abusive Clauses identified in our research, we analyze three with our ML methodology: the Illegal, the Dark, and the Gray Clauses. For these categories, we select 20 classes according to the legal source used to define them and the frequency they appear in ToS. This section describes the annotated clauses in detail, introducing their keys in the dataset, their legal sources, and their explanations in the following Table 1.
@@
lpc int Clauses that modify the liability of intermediaries of goods or services Art. 43 LPC
@@
lpc Clauses that violate a provision of the LPC that does not refer to the list of abusive clauses of Article 16. This constitutes a residual class LPC
@@
na Clauses that mention a foreign law, which is not applicable in Chile Art. 16 CC
@@
cr Clauses that include a unilateral and arbitrary modification of the ToS without the consumer being able to challenge or reject the change Art. 16 a) LPC
ter Clauses that include a unilateral and arbitrary termination of services Art. 16 a) LPC
@@
des risk Clauses that impose risks or costs on the consumer that are unrelated to the provider arising from external factors to the contractual relationship, such as force majeure Art. 16 g) LPC
des def Clauses that impose obligations on the consumer to defend, indemnify, or cooperate in a potential litigation in favor of the provider Art. 16 g) LPC
@@
bfe Clauses that go against the requirements of good faith taking into consideration the purpose of the contract. This constitutes a residual class Art. 16 g) LPC
Table 1: The annotation scheme consists of the potentially abusive clauses’ keys in the dataset, their legal sources, and their explanations
@@ 3.7 Corpus
The 50 ToS analyzed comprise a total of 6,523 clauses. If there were multiple versions of the ToS available, we selected the most recent one for Chilean customers.
The annotations were performed by two Chilean legal experts [^5] A licensed lawyer and professor with PhD in Commercial Law and a graduate with a Bachelor of Law undergoing mandatory state practice to obtain their professional license from the Chilean Supreme Court., who we recruited in-house. We used a three-step process to ensure quality control of the annotation process by expert review [43]. The first expert fully annotated a contract, and then the second expert reviewed and commented on the annotations. Ultimately, both experts discussed and agreed on the final labels. Due to the expert review, we do not report agreement measures such as Cohen’s kappa [43].
The experts first segmented the text along its paragraphs into clauses. Then, they applied the annotation scheme on the clauses and marked them with one or more keys, refer to Table 1 for details. Of the 20% that are identified as abusive, around 45% are categorized as illegal, 26% as dark, and 28% as gray. For a detailed breakdown, refer to Table 8. Following Section 3.6, we aggregate and recombine the annotated clauses into the datasets Illegal, Dark, and Gray Clauses. For each category, we define the binary classification task of detection and the multi-label multi-class task of classification. We eliminate any clause that comprises fewer than seven words. This yields a total corpus size of 5,209 annotated clauses that we use to construct six datasets.
For each detection task, we select one of the three categories of potentially abusive clauses and reduce its labels to the uniform label "abusive". Then, we combine it with the non-abusive clauses, see Table. 2 for an overview. This pre-processing step emphasizes each category’s unique challenges and creates isolated detection tasks. This is especially interesting for the novel category of Illegal Clauses, which has not been studied previously. For future work, we envision creating one model that detects the combined categories at once.
The classification task then consists of only the paragraphs that are labeled potentially abusive, see Table. 3 for an overview. These tasks are inherently multi-label, and it is common to find combinations of labels for a single clause, see the co-occurrence in Fig. 1. However, our corpus exhibits varying levels of co-occurrence, ranging from the almost pure annotations in Illegal Classification to the high rates of co-occurrence in Dark Classification and particularly in Gray Classification.
@@
Since some combinations of multi-label annotations are extremely rare, we employed an iterative stratified sampling method [59] to account for the significant disparity in the number of distinct label sets. This method ensures a highly representative split into training, validation, and testing, given the complexity of the annotations. Our use of this method involves dividing each subset into training and testing sets by a ratio of 70% to 30%, and then further subdividing the training set into training and validation sets, with a ratio of 70% to 30%.
The iterative stratified sampling technique [59] for the multi-label data causes the total sum of instances to be different to that of the corresponding detection task to sample a balanced split. For example, the algorithm selectively over-samples to stratify the Illegal Classification dataset split to a total of 643 samples instead of 603 in the detection task.
@@ 4 Methodology
We employ two learning paradigms to detect and classify potentially abusive clauses. First, we fine-tune transformer-based models. Second, we evaluate the few-shot performance of commercial and open LLMs. We follow [13] and include a linear Support Vector Machine [16] as a baseline. This SVM uses TF-IDF features for the top-$K$ frequent $n$-grams ($n\in[1,2,3]$).
@@ 4.1 Fine-Tuning Classifier Models
  • BERT [18] is an encoder model pre-trained on a corpus of English texts. It performs masked language modeling with a vocabulary of 32K tokens and up to a sequence length of 512 tokens. Longer texts are usually truncated. We include it as an essential baseline method.
  • Multilingual BERT (M-BERT) [18] is trained in 104 languages, but otherwise, it’s the same as BERT. We include it in our study because M-BERT generates multilingual representations that are easy to fine-tune and achieve excellent performance in Spanish-language tasks [51].
@@
For the detection tasks, we use a binary cross-entropy loss. For the multi-label classification task with $n$ classes, we use the binary cross entropy loss with logits, where a binary vector of length $n$ encodes the possible classes. We report the micro-F1 and macro-F1 scores for all models and datasets since the classes are highly imbalanced. Besides these, we report precision and recall metrics for detailed inspection. To calculate the multi-label metrics, we convert the networks’ sigmoid output signal to binary predictions for the $n$ classes [13].
@@ 4.2 Prompting Large Language Models
The evaluation of the few-shot capabilities of Large Language Models focuses on models that explicitly support multilingual data and have sufficiently long context lengths to process the constructed long few-shot queries. This excludes models that can only process fewer than 8,000 tokens at a time. We categorize the selected LLMs into commercial or open models. For an overview, refer to Table 4.
@@
We first based our prompt on a template that was initially proposed by LegalBench [29]. However, preliminary experimentation showed limited generalization with the different models. Errors included hallucinated class labels, extremely long answers, and high variability in prompt length in higher few-shot counts, leading to slow responses. Hence, we adapted the prompt in two ways to mitigate some of the problems: i) we prompt more concise outputs, ii) and we sample more representative few-shot instances. More advanced prompting techniques, such as automated reasoning or the chain-of-thought method [58] or the incorporation of structured knowledge via retrieval-augmented generation [15], as well as the qualitative evaluation of model reasoning, are considered as interesting future work.
We describe our prompt in the following. A header describes the task and provides a list with possible labels to choose from:
Given the following examples of clauses and their labels, predict the last clause by their label. Respond with the correct labels only. Do not explain your decision. Here are all possible options, followed by typical examples. [^6] We use the Spanish translation: ”Dados los siguientes ejemplos de cláusulas chilenas y sus etiquetas, prediga la última cláusula según su etiqueta. Responda sólo con las etiquetas correctas. No explique su decisión. A continuación, se indican todas las opciones posibles, seguidas de ejemplos típicos.”
This is then followed by one of the two following sentences, depending on whether the task is classification or detection.
A clause can have between one and three labels.
A clause can have only one label. [^7] We use the Spanish translations: ”Una cláusula puede tener entre una y tres etiquetas.” or ”Una cláusula sólo puede tener una etiqueta.”
This is followed by a list of possible clauses to predict. For instance, when classifying potentially illegal clauses, we add the following list:
@@
Label: [^8] We use the Spanish translations ”Cláusula” and ”Etiqueta”
@@ 5 Experiments
Experimental setup. We obtain pre-trained models from either the Huggingface Hub [^9] https://huggingface.co/models, transformers version 4.44.0 or the Ollama Library [^10] https://ollama.com/library, Ollama version 0.3.11. For few-shot learning, we select "instruct" variants that are tuned to follow prompts. We use 8-bit quantization, because this level of quantization increases throughput considerably, while still recommended for processing long texts because its negative effect on performance is negligible [38]. We construct the few-shot prompts from the combined training and validation datasets. For each model, we perform runs with three different random seeds and report their average test set scores with standard deviation. We fine-tune the smaller models using a binary cross entropy loss with multi-label logits and use the Adam optimizer [35] with a learning rate of 3e-5 and a batch size of 32 with up to 50 epochs. Where necessary, we apply gradient accumulation to obtain the target batch size. We regularize the training further with an early-stopping mechanism set to 10 epochs of patience. It tracks the macro-F1 score of the validation split because the datasets are imbalanced. Our fine-tuning experiments extend the code base published by LexGlue [13] [^11] https://github.com/coastalcph/lex-glue, commit 20ad76a and accelerate training using mixed precision on NVIDIA Ada Lovelace GPUs.
Since the datasets are highly imbalanced we do not only report the micro-F1 score ($\mu$-F1), which provides a global measure but may be dominated by the majority classes. We also report the macro-F1 (m-F1) score, which allows us to evaluate models on the classes of each dataset equally, highlighting the performance of minority classes. For the best-performing fine-tuned models we additionally break down the results into precision and recall for each classification task, see Sec. 5.3. Our study aggregates the results from a total of 426 experiments.
@@ 5.1 Detection
We demonstrate the ability of LLMs to detect potentially abusive clauses in the Chilean ToS corpus. To this end, we conducted large-scale experiments with a diverse set of models on the three detection datasets of the corpus and reported the $\mu$-F1 and m-F1 scores and their standard deviation per model over three runs. The results in Tab. 5 are subdivided into four groups of models: i) an SVM baseline that was trained from scratch, ii) six transformer models pre-trained on English, iii) eight transformer models pre-trained with Spanish or multiple languages, and iv) four LLMs queried via few-shot learning.
@@
Table 5: Experimental results for the three detection datasets. For convenience, we print the overall best results in bold and underline the results that are best per group. The results show that detection is feasible. Moreover, they indicate a difficulty-ranking of the corpora with the Illegal Detection dataset on the one hand being easier to predict and Dark Detection and Gray Detection on the other hand harder, resulting in lower detection rates.
@@
On the Dark Detection dataset, the baseline SVM delivers the highest $\mu$-F1 Score of 96% and BETO uncased the highest m-F1 Score of 80%. The Dark Clauses are harder to detect than Illegal Clauses. This may be due to their relatively lower occurrence in the corpus and their imbalance, leading to lower m-F1 scores. Most English transformers exhibit a very high variance in m-F1 scores and significant drops in performance. The regularization may not completely eliminate overfitting to more frequent classes, possibly leading to this problematic higher variance. However, more appropriate pre-training remedies these problems as multilingual or Spanish-language pre-training demonstrates. RoBERTalex performs on par with M-BERT cased and only falls behind BETO and M-BERT uncased by 2% m-F1 Score. Finally, few-shot learning is ineffective and the best models are behind by 10% in $\mu$-F1 and 14% $m$-F1.
@@
In summary, our proposed detection datasets present their unique challenges and our analysis emphasizes the importance of selecting a suitable model architecture and dataset characteristics, such as pre-training corpora. However, the strong performance of the SVM and small, appropriately pre-trained transformers demonstrates, that the detection of potentially abusive clauses in the Chilean ToS corpus is possible, with the rates of $96\%$ $\mu$-F1 and $89\%$ m-F1 on Illegal Detection, $96\%$ $\mu$-F1 and $80\%$ m-F1 on Dark Detection, $96\%$ $\mu$-F1 and $79\%$ m-F1 on Gray Detection. This allows us to separate the detection from subsequent classification, which we will discuss in the next section.